# Efficient Data Selection for Training Genomic Perturbation Models

### v5

@@ Abstract
Genomic studies, including CRISPR-based Perturb-seq analyses, face a vast hypothesis space, while gene perturbations remain costly and time-consuming. Gene perturbation models based on graph neural networks are trained to predict the outcomes of gene perturbations to facilitate such experiments. Due to the cost of genomic experiments, active learning is often employed to train these models, alternating between wet-lab experiments and model updates. However, the operational constraints of the wet-lab and the iterative nature of active learning significantly increase the total training time. Furthermore, the inherent sensitivity to model initialization can lead to markedly different sets of gene perturbations across runs, which undermines the reproducibility, interpretability, and reusability of the method. To this end, we propose a graph-based data filtering method that, unlike active learning, selects the gene perturbations in one shot and in a model-free manner. The method optimizes a criterion that maximizes the supervision signal from the graph neural network to enhance generalization. The criterion is defined over the input graph and is optimized with submodular maximization. We compare it empirically to active learning, and the results demonstrate that despite yielding months of acceleration, it also improves the stability of the selected perturbation experiments while achieving comparable test error.
@@ Introduction
Genomic research enables the study of genetic factors underlying various health conditions, opening new avenues for therapeutic development. Techniques such as CRISPR interference (CRISPRi) and PerturbSeq barrangou2016applications have revolutionized the landscape of genomic experimentation by enabling high-throughput screenings. However, the majority of CRISPR-based PerturbSeq experiments are restricted to hundreds of single gene perturbations dixit2016perturb ; adamson2016multiplexed ; peidli2022scperturb due to budget and time constraints, despite having over 20,000 potential gene targets. To assist exploring this vast space, machine learning models have been developed to predict the outcome of gene perturbations on a given cell roohani2024predicting ; bereket2024modelling ; bertin2023recover ; gaudelet2024season ; lopez2023learning ; lotfollahi2023predicting ; geiger2023systematically .
@@
Recent state-of-the-art methods often leverage graph-based models that use gene-gene interaction networks for perturbation prediction roohani2024predicting . These require a substantial number of training samples that stem from genetic experiments pertaining to a specific hypothesis, rendering training costly and time-consuming because it requires feedback from a wet-lab. Active learning has been proposed to address this problem huang2024sequential by defining an iterative interaction between the wet lab and the base model to run experiments for genes that will optimize the training, as shown in Fig. 1(a). This ensures that resources are not wasted in non-informative genetic experiments. However, in this setting, the training can take months, because the process typically has a few iterations ($\leq 5$huang2024sequential and each iteration translates to a CRISPRi-based PerturbSeq experiment that may take 3-5 weeks gasperini2019genome , without accounting for operational delays due to communication or computational hurdles like model retraining.
Another problem with active learning is that it relies on the model to decide the next batch of genes to train on settles2009active . However, the model is initialized randomly, and the first gene selection is by default random, leading to significantly different training sets whenever the model is retrained. This instability introduces several practical challenges. First, the resulting gene sets are less reusable: if different runs select different genes, the collected data may be too model-specific to support downstream analyses and integration with other studies. Second, thrustworthiness is undermined—if selected perturbations vary with initialization, it becomes unclear whether they reflect a true biological signal or model noise. Third, reproducibility suffers, as the same training strategy yields divergent selections in independent experiments. To mitigate this risk, we need methodologies that disentangle the model with the selection strategy. Although density-based active learning is model-independent hacohen2022active , it is not applicable in our context because the genes are not represented numerically (e.g., with features, etc.). Their embeddings are learned through the model roohani2024predicting and active learning acts upon that space huang2024sequential , which renders these selections also model-based.
In the face of these challenges, in this work, we develop a training strategy for gene expression models that is i) significantly faster, ii) operationally simpler, iii) more stable. To this end, we explore data filtering methods kirsch2025active , which select samples before training the model. Such an approach has several advantages that support real-world adaptation:
  • Decrease experimental time. Data filtering selects the full batch of perturbations upfront instead of iteratively. Since Perturb-seq assays are inherently parallelizable, this approach can complete in the time required for a single active learning cycle—resulting in months of real-world acceleration.
  • Reduced operational complexity. Active learning requires repeated coordination between the machine learning and wet-lab teams, model retraining, and acquisition computations. These steps introduce computational, logistic, and communication delays as well as potential errors, which compound across cycles and are diminished with a one-shot data filtering strategy.
  • Independence of model initialization. Since data filtering does not rely on the model, it produces a consistent set of selected genes across different runs. This stability improves reproducibility, trustworthiness, and reusability of the resulting data in downstream analyses.
@@
(b) Data filtering.
Figure 1: The difference between active learning and data filtering for training GEARS roohani2024predicting .
The only search space we can use to define data filtering is the knowledge graph, which is defined between genes based on prior knowledge and is used for message passing by the model. To this end, we propose two graph-based methods to build the train set for one of the state-of-the-art models, GEARS roohani2024predicting , although our methods can be combined with any graph-based predictor. We rely on a theorem for GNN generalization to define a proxy criterion to guide the sample selection based on the input knowledge graph, as seen in Fig. 1(b). The criterion is proven submodular, and it is optimized greedily to build the training set. Our contributions are as follows:
  • Two novel data-filtering criteria (GraphReach and MaxSpec), that select genes from the knowledge graph such that GEARS generalizes better compared to random selection. GraphReach is optimized using a submodular maximization algorithm, and MaxSpec using a greedy heuristic.
  • An empirical comparison including two active learning methods (one of which is state of the art in our problem), indicating that GraphReach achieves at least 5× faster training and increased stability without essentially sacrificing accuracy.
@@
The body of related work can be broadly divided into two categories. The first concerns optimal experimental design, where the goal is to select gene perturbations that maximize a downstream causal effect (e.g., a phenotype). The objective is to choose a subset of perturbations (treatments) from a large action space in order to maximize an expected outcome variable, such as T-cell activation  mehrjou2021genedisco . Methods such as Bayesian optimization and online learning have been applied in this context pacchiano2022neural ; lyle2023discobax ; pacchiano2022neural , and are evaluated based on the number of high-reward interventions discovered.
@@
The second branch of relative literature focuses on efficient training strategies for models that predict the gene expression profile following perturbation in single cells. In this setting, the aim is not to optimize a causal phenotype but to train a model that generalizes well to unseen perturbations. This is particularly useful in Perturb-seq experiments, which measure the transcriptomic effects of perturbations via single-cell RNA sequencing. Due to the high cost and long duration of each experimental cycle, active learning methods have been proposed to efficiently select the most informative perturbations for training huang2024sequential .
The closest method to our approach is IterPert huang2024sequential , which uses active learning and prior multimodal knowledge to build a train set for GEARS. Since each iteration can take 3-5 weeks in the wet-lab, the number of iterations is diminished to 5 in contrast to  50 in optimal experimental design mehrjou2021genedisco . The problem is addressed from the perspective of active learning under budget hacohen2022active with the inclusion of prior imaging and perturb-seq studies. In fact, prior multimodal data was so effective, that it produced state of the art results without active learning, i.e., the model IterPert-prior-only. However, such data is not commonly available.
@@
Our methods differ from prior work in several important ways. First, we propose graph-based data filtering approaches that select all perturbations prior to model training, thereby requiring only a single experimental round. Since Perturb-seq experiments can be parallelized, this reduces the experimental duration from 5-6 months to approximately one month. Second, unlike IterPert, our approach relies only on an open-access ontology, making it widely applicable across settings where curated multimodal data is unavailable. Third, by removing model-driven selection, the methods achieve stable and reproducible gene selection. This improves the reusability and interpretability of the resulting data. Finally, one of the proposed methods achieves competitive test error compared to the state-of-the-art active learning method that does not use prior multimodal data, TypiClust huang2024sequential .
@@
In this section, we first formulate the problem, clarify the theoretical motivation for our approaches and we introduce the two methods for data filtering.
@@
The model parameters $\theta^{S}$, trained only on the selected set $S$, represent a graph neural network (GNN) such as GEARS. The loss function is the mean squared error between the predicted and the observed expression:
@@
Data Filtering to Maximize Supervision
We now provide relevant literature and an argument based on gradient flow from our train set to motivate the guiding principle of our methods, which is to maximize the set of reachable nodes from our chosen train set.
Since we refrain from utilizing the model’s predictions, we turn to methodology that resembles traditional density-based active learning, such as Core-sets sener2017active , which aims to cover as much of the input space as possible. Core-sets rely on the fact that if we assume each training sample to cover a ball with radius $\delta$ around them, then choosing the training set to maximize coverage of the dataset reduces the generalization error. We also know from the literature that labeling a node results in an information gain for every other node in their receptive field, i.e., $k$-hop neighborhood wang2020unifying ; zhang2022information . Finally, we know that the generalization is improved as we reduce the geodesic distance between the training and the test nodes ma2021subgroup . These findings support our intuition that covering a larger part of the graph with training nodes should be beneficial for the model.
@@
where $\mathbf{D}$ is the diagonal degree matrix, $\mathbf{W}_{0}\in\mathbb{R}^{N\times d}$ is the embedding lookup table, $\hat{\mathbf{A}}$ is the normalized adjacency with self-loops to the power of $k=1$ (which is the default parameter chosen in the GEARS model), $\mathbf{W}_{1}\in\mathbb{R}^{d\times d^{\prime}}$ are the SGC’s weights, and $\mathbf{O}\in\mathbb{R}^{N\times N}$ is a row-wise one-hot encoding of the candidate gene perturbations. We base our methodology on the fact that message-passing GNNs update only the representations in $\mathbf{W}_{0}$ of nodes that are within the receptive field of the supervised nodes. To formalize this, consider the gradient of the loss $\mathcal{L}$ with respect to the identity embeddings $\mathbf{W}_{0}$ as a function of $\frac{\partial\mathcal{L}}{\partial\mathbf{H}}$:
Thus, the full gradient becomes:
which shows that the gradient of $\mathbf{W}_{0}[i]$ depends only on the gradients of its neighbors in $\hat{\mathbf{A}}$. However, $\frac{\partial\mathcal{L}}{\partial\mathbf{H}[j]}=0$ if $j\notin S$. Hence the supervision signal is propagated solely to nodes reachable from the train set, meaning that the identity embedding $\mathbf{W}_{0}$ that is learned for each gene, is updated only if it lies within the receptive field of supervised nodes.
This means that if we choose the train set such that we maximize the reachable nodes, we will maximize in tandem the number of $\mathbf{W}_{0}$’s adjusted to the supervision signal. This is beneficial because, since GEARS is inherently semi-supervised, expanding the supervision allows for the representations of test samples or their neighbors to get updated, which leads to better generalization. This clarifies how covering larger parts of the graph can potentially improve the prediction. Therefore, to maximize the reach of the supervision signal, we should aim to maximize the number of nodes reached by the training set $S$.
Figure 2: An overview of the data filtering methodology. Our sole input is the knowledge graph from GO roohani2024predicting and a list of candidate perturbations. The data filtering algorithm, such as GraphReach selects the set of gene perturbations, and they are given to the wet lab for the experimental part. Finally, the single-cell gene expression data is given to GEARS for training and validation.
@@
We define our data filtering criterion as a function that maximizes the number of nodes reached by the node set $S$, i.e, the receptive field. Specifically for a new node $v$, the criterion is defined as the additional reachability that $v$ adds in the set of train perturbations $S$
@@
The subscript $L$ is omitted in subsequent discussion as it remains stable throughout all experiments. The criterion function selects greedily the node $v$ maximizing $\alpha(v,S)$ in every iteration. This is fundamentally suboptimal baykal2021low ; kirsch2019batchbald , but like many active learning methods lyle2023discobax ; pacchiano2022neural ; tigas2022interventions ; sussex2021near , we will prove our criterion is monotonic and submodular to achieve a guarantee that the result will be at least 11/e1-1/e close to the optimal nemhauser1978analysis .
Lemma 1.
@@
The reader is referred to the Appendix for the whole proof. Let two sets $S_{t}\subseteq S_{t+1}\subseteq V$ and any node $v\notin S_{t+1}$. Since $S_{t}\subseteq S_{t+1}$, every node reachable from $S_{t}$ is also reachable from $S_{t+1}$, meaning $R(S_{t})\subseteq R(S_{t+1})\Rightarrow~|R(S_{t})|\leq|R(S_{t+1})|$, which proves monotonicity. By definition, adding a node earlier ($\alpha(v,S_{t})$) adds at least as many reachable nodes as adding it later ($\alpha(v,S_{t+1})$), since the nodes in the cut of $S_{t}$ and $S_{t+1}$ can belong to $R(v)$ as well: $R(v)\cap(R(S_{t+1})\setminus R(S_{t}))\geq 0$. This means that any node that appears in $R(v)$ and in $(R(S_{t+1})\setminus R(S_{t}))$ was new for $R(S_{t})$ but is not new for $R(S_{t+1})$. Hence, the additional reachability that $v$ can bring to $S_{t+1}$ compared to the one it can bring to $S_{t}$ is diminished. This sketches out the proof that the reachability is submodular.
The greedy approach requires testing every available perturbation for every addition during a cycle, i.e., $\mathcal{O}(NT)$ where $T$ is the total training budget in samples. We employ a cost effective lazy forward approach to accelerate the acquisition step without loss of accuracy leskovec2007cost . The final data filtering algorithm, called GraphReach from graph reachability, is shown in Alg. 1, and an overview of the overall step-by-step data filtering approach can be seen in Fig. 2.
@@
MaxSpec
Another line of research suggests that GNN generalization improves if the largest singular value (i.e., spectral norm) of the message passing operator is diminishing ju2023generalization . We can translate this as choosing a perturbation such that the underlying graph formed by the nodes that are reachable from $S$, has the smallest possible spectral norm of the respective diffusion matrix. The diffusion matrix is $\mathbf{D}^{-1/2}(\mathbf{A}_{S}+\mathbf{I}d)\mathbf{D}^{-1/2}$ with $\mathbf{A}_{S}=A[R(S),R(S)]$ is the subgraph induced by $R(S)$. Hence we define a new criterion function similar to Eq. 10, but instead of reachability, we utilize the spectral norm as:
$M(\mathbf{A}_{S})=\|\mathbf{D}^{-1/2}(\mathbf{A}_{S}+\mathbf{I}d)\mathbf{D}^{-1/2}\|_{2}$ (12)
We still rely on a cost-effective lazy forward method, hence the algorithm MaxSpec (from maximizing spectral norm) is the same Alg. 1, but substituting function $R$ with $M$.
@@
As mentioned in the introduction, data filtering is significantly faster than active learning due to high-throughput genomics platforms like Perturb-seq being explicitly designed for parallelized experiments. This translates to a 5-fold acceleration in our context. Besides speed, here we quantify the changes in other dimensions of the problem and visualize the tradeoff. Specifically, we address these questions:
@@
  • Test Error: How is the test error of the model affected by the training procedure?
  • Tradeoff: Does the overall tradeoff justify data filtering approaches?
To this end, in this section, we first describe the data used for the experiments, including the gene perturbation datasets and the knowledge graph, the benchmarks used for comparison, as well as the experimental design and the evaluation methods. Afterwards, we showcase the performance of the methods and analyze them to derive conclusions about the soundness of the proposed techniques. The code of the experiments, along with details on computing infrastructure is provided in the supplementary material.
@@
We test our methods in two genomic datasets stemming from PerturbSeq experiments, single-cell CRISPR-based screenings on K562 chronic myeloid leukemia cells:
  • Adamson: A study that investigates the mammalian unfolded protein response of K562 cells to CRISPR interference such as repression. After data curation, it contains 81 binary gene perturbations with approximately 800 cells each adamson2016multiplexed .
  • Replogle: A study that systematically maps genotype-phenotype relationships using CRISPR knockouts. After data curation, it includes 1087 binary gene perturbations with approximately 150 cells each replogle2022mapping .
The datasets are diverse in terms of supervision (150 to 800 cells) and number of perturbations (1087 to 82). We chose them in order to juxtapose the method performances in these two different settings and identify which methods perform adequately in both, taking into consideration the scarcity of relevant datasets.
The graph we focus on for both datasets is based on pathway information from GO51 gene2004gene . Each gene is associated with a number of pathway GO51 terms. The Jaccard similarity between the sets of pathways of two genes, i.e., the fraction of shared pathways, is used to calculate the strength of the edge between them. The graph is sparsified by keeping a predefined number of the most important neighbors for each node. The final graph contains over 9,800 nodes and 200,000 edges, exactly as in roohani2024predicting . It should be noted that GEARS utilizes the whole knowledge graph to learn representations and uses the gene perturbation datasets for supervision only in a semi-supervised manner. The data filtering methods follow suit and utilize the whole graph. For example, the reachability of a gene perturbation in Replogle is not defined only on the 1088 candidate perturbations, but on the whole 9,800 nodes.
(a) Adamson.
(b) Replogle.
Figure 3: The total number of unique genes perturbed in active learning experiments with 3 different random seed initializations. The variation in the data filtering methods is induced solely by different $k$-fold splits. The difference with the other methods corresponds to the sensitivity of the corresponding gene selection process to the model initialization and the first gene selection, which are both random.
@@
The benchmark models are:
  • Random selection from the available gene perturbations. This is the prevalent practice in application because of its speed and simplicity.
  • ACS-FW pinsler2019bayesian is a Bayesian batch active learning model. It selects perturbations such that the new posterior distribution of the model’s parameters approximates the expected posterior when the whole dataset is available. We utilized the version from the BMDALreg library holzmuller2023framework , which was one of the top-performing methods in similar experiments on active learning for GEARS huang2024sequential .
  • TypiClust hacohen2022active is the state-of-the-art active learning algorithm on our problem that does not rely on multimodal priors huang2024sequential . It has outperformed a plethora (8 to be precise) of models.The algorithm clusters the candidate perturbations based on the final graph layer representation from GEARS. Within each cluster, the typicality is quantified as the inverse of the average distance between each sample and an example’s $K$-nearest neighbors, with $K=20$. The most typical sample is selected.
Note that IterPert does not operate without multi-modal data and hence can not be utilized in our setting.
(a) Adamson.
(b) Replogle.
Figure 4: Tradeoff between training time, predictive error (MSE), and stability of gene selection (the dot size is analogous to mean Jaccard similarity across seeds). Each point corresponds to a different perturbation selection strategy. GraphReach achieves low error and high stability in a single experimental cycle (1 month), compared to model-dependent active learning methods (e.g., TypiClust), which require five cycles and exhibit less stable gene selection.

### v6

@@ Abstract
Genomic studies face a vast hypothesis space, while interventions such as gene perturbations remain costly and time-consuming. To accelerate such experiments, gene perturbation models predict the transcriptional outcome of interventions. Since constructing the training set is challenging, active learning is often employed in a “lab-in-the-loop” process. While this strategy makes training more targeted, it is substantially slower, as it fails to exploit the inherent parallelizability of Perturb-seq experiments. Here, we focus on graph neural network–based gene perturbation models and propose a subset selection method that, unlike active learning, selects the training perturbations in one shot. Our method chooses the interventions that maximize the propagation of the supervision signal to the model. The selection criterion is defined over the input knowledge graph and is optimized with submodular maximization, ensuring a near-optimal guarantee. Experimental results across multiple datasets show that, in addition to providing months of acceleration compared to active learning, the method improves the stability of perturbation choices while maintaining competitive predictive accuracy.
@@ Introduction
Genomic research enables the study of genetic factors underlying various health conditions, opening new avenues for therapeutic development. Techniques such as CRISPR interference and PerturbSeq (barrangou2016applications, ) have revolutionized the landscape of genomic experimentation by enabling high-throughput screenings. However, the majority of CRISPR-based PerturbSeq experiments are restricted to hundreds of single gene perturbations (dixit2016perturb, ; adamson2016multiplexed, ; peidli2022scperturb, ) due to budget and time constraints, despite having over 20,000 potential gene targets. To assist exploring this vast space, machine learning models have been developed to predict the outcome of gene perturbations on a given cell (bereket2024modelling, ; gaudelet2024season, ; lopez2023learning, ; lotfollahi2023predicting, ).
@@
Recent methods often leverage graph-based models that use gene-gene interaction networks for perturbation prediction (roohani2024predicting, ). They require a substantial number of training samples that stem from genomic experiments pertaining to a specific hypothesis, rendering training costly and time-consuming because it requires feedback from a wet-lab. Active learning has been proposed to address this problem (huang2024sequential, ) by defining an iterative interaction between the wet lab and the base model to run experiments for genes that will optimize the training, as shown in Fig. 1(a). This ensures that resources are not wasted in non-informative genomic experiments. However, in this setting, the training can take months, because despite the process having typically a few iterations ($\leq 5$) (huang2024sequential, ), each iteration translates to a CRISPRi-based PerturbSeq experiment that may take 3-5 weeks (gasperini2019genome, ; huang2024sequential, ), without accounting for operational delays due to communication or computational hurdles like model retraining.
@@
(b) Subset selection.
Figure 1: The difference between active learning and subset selection for training GEARS.
Another drawback of active learning is that it relies on a randomly initialized model to decide the first batch of genes. As a result, different runs can yield substantially different training sets, which undermines reproducibility and limits the reusability of collected data in downstream analyses. To mitigate this instability, we need selection strategies that decouple the model’s output from the training set choice. Density-based active learning is model-independent (hacohen2022active, ), but it does not apply here because genes lack predefined features; their embeddings are learned by the model itself, so active learning still operates in a model-dependent space (huang2024sequential, ).
In response, we propose a subset selection strategy that builds the training set upfront, without model input. This approach offers several practical benefits:
  • Less experimental time. Because Perturb-seq assays are parallelizable, selecting all perturbations at once allows the experiments to conclude in approximately the time of a single active learning cycle—yielding months of real-world acceleration.
  • Reduced operational complexity. It avoids repeated cycles of coordination between the wet-lab and the model, model retraining, and acquisition computations, reducing both logistical overhead and potential errors.
  • Greater stability. By decoupling the selection strategy from model initialization it ensures consistent perturbation sets across runs, improving reproducibility and reusability of collected data.
The only search space we can use to define subset selection is the knowledge graph, which is defined between genes based on prior knowledge and is used for message passing by the model. To this end, we propose a graph-based method to build the train set for one of the state-of-the-art models, GEARS (roohani2024predicting, ) as seen in in Fig. 1(b), although it can be combined with any graph-based predictor. We draw inspiration from recent findings on the relationship between train and test nodes to develop a method that selects the genes that maximize the model’s reach on the graph. We show that this selection increases the amount of genes trained, which in tandem increases the probability of having non-random embeddings in the receptive field of a test node. The criterion is proven submodular, and it is optimized greedily to build the training set with a near-optimal theoretical guarantee. Our contributions are as follows:
  • A novel subset selection method, GraphReach, that selects genes such that the supervision signal to the embeddings is maximized.
  • An empirical comparison including two active learning methods (one of which is state of the art in our problem), indicating that GraphReach achieves at least substantially faster training and increased stability without essentially sacrificing accuracy.
@@
The body of related work can be broadly divided into two categories. The first concerns optimal experimental design, where the objective is to choose a set of perturbations from a large action space in order to maximize an expected outcome variable, such as T-cell activation (mehrjou2021genedisco, ). Methods such as Bayesian optimization and online learning have been applied in this context (pacchiano2022neural, ; lyle2023discobax, ; pacchiano2022neural, ), and are evaluated based on the number of high-reward interventions discovered. These algorithms are not applicable in our case because they require a plethora of experiment rounds (e.g., 40), which are feasible with CRISPRi experiments but not when it is combined with PerturbSeq. Moreover, they focus on the experimental success of the retrieved selection rather then the efficacy of the final learning model, which typically predicts scalar values such as phenotypes not multidimensional vectors as in our problem.
@@
The second branch of relative literature focuses on efficient training strategies for models that predict the gene expression profile following perturbation in single cells. In this setting, the aim is not to optimize a causal phenotype but to train a model that generalizes well to unseen perturbations. This is particularly useful in Perturb-seq experiments, which measure the transcriptomic effects of perturbations via single-cell RNA sequencing. Due to the high cost and long duration of each experimental cycle, active learning methods have been proposed to efficiently select the most informative perturbations for training. The closest method to our approach is IterPert (huang2024sequential, ), which uses active learning and prior multimodal knowledge to build a train set for GEARS. Since each iteration can take 3-5 weeks in the wet-lab, the number of iterations is diminished to 5 in contrast to  50 in optimal experimental design (mehrjou2021genedisco, ). The problem is addressed from the perspective of active learning under budget (hacohen2022active, ) with the inclusion of prior imaging and perturb-seq studies. In fact, prior multimodal data was so effective, that it produced state of the art results without active learning, i.e., the model IterPert-prior-only. However, such data is not commonly available.
@@
Our method differs from prior work in several ways. First, we propose a method that selects the perturbations prior to model training, thereby requiring only a single experimental round to get the labels. Since Perturb-seq experiments can be parallelized, this reduces the typical duration of training the model from 5 months to approximately one month. Second, unlike IterPert, our approach relies only on an open-access ontology, making it applicable across settings where curated multimodal data is unavailable. Third, by removing model-driven selection, our methods achieve stable and reproducible gene selection. This improves the reusability and interpretability of the resulting data. Finally, the proposed method achieves competitive accuracy compared to the state-of-the-art active learning method that does not use prior multimodal data, TypiClust (huang2024sequential, ).
Figure 2: An overview of the subset selection methodology. Our sole input is the knowledge graph from GO51 (roohani2024predicting, ) and a list of candidate perturbations. The subset selection algorithm, such as GraphReach, selects the set of gene perturbations, and they are given to the wet lab for the experimental part. Finally, the single-cell gene expression data is given to GEARS for training and validation.
@@
In this section, we formulate the problem, clarify the theoretical motivation for our approaches and introduce the method for subset selection.
@@
The model parameters $\theta^{S}$, trained only on the selected set $S$, represent a graph neural network (GNN) such as GEARS (roohani2024predicting ). The loss function is the mean squared error between the predicted and the observed expression:
@@
Subset Selection to Maximize Supervision
Since we refrain from utilizing the model’s predictions, we turn to methodology that resembles traditional density-based active learning, such as Core-sets (sener2017active, ), which aims to cover as much of the input space as possible. Such methods contend that if we assume each training sample to cover all samples in a ball with radius $\delta$ around them, then choosing the training set to maximize coverage of the dataset reduces the generalization error. We also know from the literature that labeling a node results in an information gain for every other node in its receptive field, i.e., $k$-hop neighborhood (wang2020unifying, ; zhang2022information, ). Finally, we know that the generalization is improved as we reduce the geodesic distance between the training and the test nodes (ma2021subgroup, ), which is performed implicitely when the train set maximizes its reach. These findings support our intuition that covering a larger part of the graph with training nodes should be beneficial for the model.
@@
where $\mathbf{D}$ is the diagonal degree matrix, $\mathbf{W}_{0}\in\mathbb{R}^{N\times d}$ is the embedding lookup table, $\hat{\mathbf{A}}$ is the normalized adjacency with self-loops to the power of $k=1$ (which is the default parameter chosen in the GEARS model), $\mathbf{W}_{1}\in\mathbb{R}^{d\times d^{\prime}}$ are the SGC’s weights, and $\mathbf{O}\in\mathbb{R}^{N\times N}$ is a row-wise one-hot encoding of the candidate gene perturbations. We base our methodology on the fact that message-passing GNNs update only the representations in $\mathbf{W}_{0}$ of nodes that are within the receptive field of the supervised nodes. To formalize this, we now state a proposition in which we explicity calculate the gradient with respect to a particular node representation.
Proposition 1.
For the model defined in Equations 3, 4 and 5, the gradient of the loss with respect to an individual gene embedding $\mathbf{W}_{0}[i]$ is defined as
$\frac{\partial\mathcal{L}}{\partial\mathbf{W}_{0}[i]}$ $=\sum_{j=1}^{N}\hat{\mathbf{A}}[j,i]\cdot\left(\frac{\partial\mathcal{L}}{\partial\mathbf{H}[j]}\,\mathbf{W}_{1}^{\top}\right).$ (6)
Therefore, the gradient of $\mathbf{W}_{0}[i]$ depends only on the gradients of its neighbors in $\hat{\mathbf{A}}$.
The proof of Proposition 1 can be found the appendix. From Proposition 1 we can conclude that the supervision signal is propagated solely to nodes reachable from the train set, meaning that the identity embedding $\mathbf{W}_{0}$ that is learned for each gene, is updated only if it lies within the receptive field of supervised nodes.
This means that if we choose the train set such that we maximize the reachable nodes, we will maximize in tandem the number of embeddings $\mathbf{W}_{0}$ adjusted to the supervision signal. This is beneficial because, since GEARS is inherently semi-supervised, expanding the supervision allows for the representations of test samples or their neighbors to get updated. Having non-random embeddings in the receptive field of the test nodes should in principle lead to better generalization. It should be noted that while our analyzes relies on the architecture of GEARS, we expect the conclusion to generalize to a number of similar semi-supervised GNNs. This clarifies how covering larger parts of the graph can potentially improve the prediction. Therefore, to maximize the reach of the supervision signal, we should aim to maximize the number of nodes reached by the training set $S$.
@@
We define our subset selection criterion as a function that maximizes the number of nodes reached by the node set $S$, i.e, the receptive field. Specifically for a new node $v$, the criterion is defined as the additional reachability that $v$ adds in the set of train perturbations $S$
@@
The subscript $L$ is omitted in subsequent discussion as it remains stable throughout all experiments. The criterion function selects greedily the node $v$ maximizing $\alpha(v,S)$ in every iteration. This is fundamentally suboptimal (baykal2021low, ; kirsch2019batchbald, ) but like many active learning methods (lyle2023discobax, ; pacchiano2022neural, ; tigas2022interventions, ; sussex2021near, ), we will prove our criterion is monotonic and submodular to achieve a guarantee that the result will be at least 11/e1-1/e close to the optimal (nemhauser1978analysis, ).
Proposition 2.
@@
The reader is referred to the Appendix for the whole proof. Consider two sets $S_{t}\subseteq S_{t+1}\subseteq V$ and any node $v\notin S_{t+1}$. Since $S_{t}\subseteq S_{t+1}$, every node reachable from $S_{t}$ is also reachable from $S_{t+1}$, meaning $R(S_{t})\subseteq R(S_{t+1})\Rightarrow~|R(S_{t})|\leq|R(S_{t+1})|$, which proves monotonicity. By definition, adding a node earlier ($\alpha(v,S_{t})$) adds at least as many reachable nodes as adding it later ($\alpha(v,S_{t+1})$), since the nodes in the cut of $S_{t}$ and $S_{t+1}$ can belong to $R(v)$ as well: $R(v)\cap(R(S_{t+1})\setminus R(S_{t}))\geq 0$. This means that any node that appears in $R(v)$ and in $(R(S_{t+1})\setminus R(S_{t}))$ was new for $R(S_{t})$ but is not new for $R(S_{t+1})$. Hence, the additional reachability that $v$ can bring to $S_{t+1}$ compared to the one it can bring to $S_{t}$ is diminished. This sketches out the proof that the reachability is submodular.
Despite the theoretical guarantee, the greedy approach to maximizing reachability over the graph requires still requires testing every available perturbation for every addition during a cycle, i.e., $\mathcal{O}(NT)$ where $T$ is the total training budget in samples. This can be substantial depending on the graph or the experiment’s scope. To this end, we employ a cost effective lazy forward approach to accelerate the acquisition step without loss of accuracy (leskovec2007cost, ). The final subset selection algorithm, called GraphReach from graph reachability, is shown in Alg. 1, and an overview of the overall step-by-step subset selection approach can be seen in Fig. 2. Finally, it should be noted that the method is easily extended to combinatorial perturbations. The set $S$ includes the candidate sets of perturbations, and the reachability $R(S)$ is the total number of unique nodes reached by the joint candidate node set. If two subsets of $S$ overlap, then the joint set ensures the individual gene appears only once.
@@
As mentioned in the introduction, subset-selection is significantly faster than active learning due to high-throughput genomics platforms like Perturb-seq being explicitly designed for parallelized experiments. This translates to a many-fold acceleration in our context. Besides speed, here we quantify the changes in other dimensions of the problem, addressing these questions:
  • Accuracy: How is the generalization of the model affected by the training procedure?
@@
  • Robustness: Is noise in the knowledge graph able to erode the accuracy of the model?
To this end, in this section, we first describe the data used for the experiments, including the gene perturbation datasets and the knowledge graph, the benchmarks used for comparison, as well as the experimental design and the evaluation methods. Afterwards, we showcase the performance of the methods and analyze them to derive conclusions about the soundness of the proposed techniques. The code of the experiments, along with details on computing infrastructure, is provided in the supplementary material.
@@
We test our methods in four single-cell genomic datasets stemming from PerturbSeq experiments, following the literature (roohani2024predicting, ) as seen in Tab. 1. The datasets are diverse in terms of the number of perturbations and the number of samples per perturbation. The Norman dataset includes combinatorial perturbations (up to 2 genes) and the rest contain data on single perturbations.
Table 1: The number of distinct gene perturbations in each datasets (single or combinatorial) along with the average number of cells (samples) per perturbation.
Dataset Perturbation Number Average Cell Count
Adamson 81 800
K562 1,087 150
RPE1 1,535 105
Norman 277 322
Akin to the literature, the graph is based on pathway information from GO51 (gene2004gene, ). Each gene is associated with a number of pathway GO51 terms. The Jaccard similarity between the sets of pathways of two genes, i.e., the fraction of shared pathways, is used to calculate the strength of the edge between them. The graph is sparsified by keeping a predefined number of the most important neighbors for each node. The final graph contains over 9,800 nodes and 200,000 edges, exactly as in (roohani2024predicting, ). It should be noted that GEARS utilizes the whole knowledge graph to learn representations and uses the gene perturbation datasets for supervision only in a semi-supervised manner. GraphReach follows suit and utilizes the whole graph for selection.
@@
As stated in Section Related Work, there is currently still a lack of methods addressing our tackled problem. IterPert does not operate without multi-modal data and hence cannot be utilized in our setting. Thus, we rely on these benchmarks:
  • Baseline represents the random selection from the available gene perturbations. This is the prevalent practice because of its speed and simplicity. It represents the vanilla usage of GEARS.
  • ACS-FW (pinsler2019bayesian, ) is a Bayesian batch active learning model. It selects perturbations such that the new posterior distribution of the model’s parameters approximates the expected posterior when the whole dataset is available. We utilized the version from the BMDALreg library (holzmuller2023framework, ), which was one of the top-performing methods in similar experiments (huang2024sequential, ).
  • TypiClust (hacohen2022active, ) is the state-of-the-art active learning method on this problem as it has outperformed 8 active learning models (huang2024sequential, ). The algorithm clusters the candidate perturbations based on the final graph layer representation from GEARS. Within each cluster, the typicality is quantified as the inverse of the average distance between each sample and an example’s $K$-nearest neighbors, with $K=20$. The most typical sample is selected.
Figure 3: Accuracy in the test set with confidence intervals compared to the total physical time required for the training due to wet-lab iterations. GraphReach and Baseline do not require model-input hence they are trained with only one cycle of wet lab experiments taking 30 days. TypiClust and ACS-FW inherently need numerous cycles, and they are each run for 90 and 150 days to highlight the role that the number of cycles plays in the accuracy.